OpenStudy (anonymous):

The first five terms of a sequence are shown below. 4, 7, 10, 13, 16 If the nth term of this sequence is represented by f(n), which of the following functions best represents this sequence? f(n) = 3n - 1; n ≥ 1 f(n) = 3n + 1; n ≥ 1 f(n) = 4n - 1; n ≥ 1 f(n) = 4n + 1; n ≥ 1

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

This is an arithmetic progression, with the first term 5 and common difference: 4. The nth term of the sequence: a, a + d, a + 2d, a + 3d, ..... is: a + (n-1)d

OpenStudy (gorv):

seee in each term multiplied by 3and 1 is added

OpenStudy (gorv):

first term 3*1+1=4

OpenStudy (rianiscool):

i think so to

OpenStudy (gorv):

second 3*2+1=7

OpenStudy (gorv):

3*3+1=10

OpenStudy (anonymous):

The first term is a. Here a = 5. The common difference is d. Here d = 4. Put it in the formula for the nth term and simplify

OpenStudy (gorv):

so it will be 3n+1

OpenStudy (anonymous):

It is simple to verify too. Just put n = 1, 2, 3, .... in 4n + 1 and see if you get each term of the sequence.

OpenStudy (anonymous):

D

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

A sequence has its first term equal to 3, and each term of the sequence is obtained by adding 5 to the previous term. If f(n) represents the nth term of the sequence, which of the following recursive functions best defines this sequence? f(1) = 3 and f(n) = f(n - 1) + 5; n > 1 f(1) = 5 and f(n) = f(n - 1) + 3; n > 1 f(1) = 3 and f(n) = f(n - 1) + 5n; n > 1 f(1) = 5 and f(n) = f(n - 1) + 3n; n > 1

OpenStudy (anonymous):

For the last question? @woooooooooowwwww

OpenStudy (anonymous):

oh what?

OpenStudy (gorv):

f(1)=3 f(n)= previous term +5 previous term = f(n-1) f(n)=f(n-1)+5

OpenStudy (anonymous):

???????

OpenStudy (anonymous):

so it would be.......